OpenStudy (anonymous):

Please help!!! Formula is A > (wh)+ 0.5(bh). solve for b

OpenStudy (anonymous):

@Compassionate @campbell_st

OpenStudy (anonymous):

@Jamierox4ev3r

OpenStudy (anonymous):

I don't understand it at all, how to do it

HanAkoSolo (jamierox4ev3r):

well what you do is solve for the specific variable on its own. Do every operation possible in order to get the variable on one side. Sorry its been a while since I've done this, try asking others for help.

OpenStudy (anonymous):

ok no problem, thanks for coming

OpenStudy (anonymous):

@lalaly

HanAkoSolo (jamierox4ev3r):

no problem, here I'll do my best to work it out and if I get it, I'll let you know 'kay?

OpenStudy (anonymous):

that's fine, dont worry :) I will ask other

OpenStudy (anonymous):

@e.mccormick

OpenStudy (lalaly):

A-wh>0.5(bh) (A-wh)/0.5h >b so b< (A-wh)/0.5h

OpenStudy (anonymous):

ok you first subtracted both like this? A- wh > wh + wh 0.5(bh

OpenStudy (lalaly):

this is your question A> (wh)+0.5(bh) so i subtracted wh from both sides A-wh>0.5(bh) then divided both sides by 0.5h

OpenStudy (anonymous):

I don't really get the process solving this equation ? would mind me to show step by step?

OpenStudy (anonymous):

I m really struggling with it

OpenStudy (lalaly):

ok to make it easier.. make the "greater than" sign an "=" sign and solve it then replace the = with > when ur done so its A=(wh)+0.5(bh) first

OpenStudy (lalaly):

firs thing u subtract wh from both sides A-wh=wh + 0.5(bh) -wh so it becomes A-wh = 0.5(bh) are u following?

OpenStudy (anonymous):

yes

OpenStudy (lalaly):

ok so then u want to solve for b so u want b on one side alone now u divide both sides by 0.5h (A-wh)/0.5h = (0.5bh)/0.5h --> (A-wh)/0.5h = b now put the > back (A-wh)/0.5h >b

OpenStudy (anonymous):

thank you so much!!!!! You are explanation really helped me to understand it now :) thanks alot!!

OpenStudy (lalaly):

your welcome :)

OpenStudy (anonymous):

You are really genius !!!!!!!

OpenStudy (lalaly):

im flattered;D

OpenStudy (anonymous):

So will be it be like this b> ( A-wh)/0.5h

OpenStudy (anonymous):

because you told me place it back

OpenStudy (anonymous):

b?

OpenStudy (lalaly):

yea but when u place it back its (A-wh)/0.5h >b the sign is pointing at b when u want to rearrange the enequality it becomes b< (A-wh)/0.5h

OpenStudy (anonymous):

ok thanks a lot again for help!!!! :)

OpenStudy (lalaly):

no problem

OpenStudy (anonymous):

You are truly really amazing :)

OpenStudy (lalaly):

i'm happy to help. whenever u need help let me know

OpenStudy (anonymous):

Thank you so much for everything :)

OpenStudy (lalaly):

anytime:)